OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (anonymous):

Here's a hint: I think you have to use the slope intercept form for part of this problem. The slope intercept form is y=mx+b

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@Jadeishere

OpenStudy (jadeishere):

First step is to determine the slope of the first equation >.< so; 4x - 3y = -12 -3y = -4x - 12 y = -4/3x + 4 so the other equation has to have the same slope and pass through -3, 4.

OpenStudy (anonymous):

so I have to graph that equation and find the slope right?

OpenStudy (jadeishere):

the slope is -4/3 and the y intercept of 4x - 3y = -12 is 4

OpenStudy (jadeishere):

It'd be y = 4/3x + 8...

OpenStudy (jadeishere):

It's difficult to explain. I usually just graph the equations given and BOOM it hits me.

OpenStudy (anonymous):

lol

OpenStudy (jadeishere):

>.<

OpenStudy (anonymous):

but how do you find the equation that is parallel to it

OpenStudy (anonymous):

oh wait got it nvm mind thanks so much!

OpenStudy (jadeishere):

you graph the given options if there are any.

OpenStudy (anonymous):

I appreciate it

OpenStudy (jadeishere):

Yeah, i hope you don't mind my asking, but why'd you call me to the question?

OpenStudy (anonymous):

lol because I saw you in another chat and I thought you might be helpful and thx lol I really needed help

OpenStudy (anonymous):

btw how do I give you a medal

OpenStudy (jadeishere):

Yep :)